# Scaling Political Texts with Large Language Models: Asking a Chatbot Might Be All You Need

### v2

@@ Abstract
We use instruction-tuned Large Language Models (LLMs) such as GPT-4, MiXtral, and Llama 3 to position political texts within policy and ideological spaces. We directly ask the LLMs where a text document or its author stand on the focal policy dimension. We illustrate and validate the approach by scaling British party manifestos on the economic, social, and immigration policy dimensions; speeches from a European Parliament debate in 10 languages on the anti- to pro-subsidy dimension; Senators of the 117th US Congress based on their tweets on the left-right ideological spectrum; and tweets published by US Representatives and Senators after the training cutoff date of GPT-4. The correlation between the position estimates obtained with the best LLMs and benchmarks based on coding by experts, crowdworkers or roll call votes exceeds .90. This training-free approach also outperforms supervised classifiers trained on large amounts of data. Using instruction-tuned LLMs to scale texts in policy and ideological spaces is fast, cost-efficient, reliable, and reproducible (in the case of open LLMs) even if the texts are short and written in different languages. We conclude with cautionary notes about the need for empirical validation.
@@ 1 Introduction
Much research in the social and political sciences involves estimating the positions of actors, such as politicians or political parties, in latent ideological and policy latent spaces, such as the left-right or liberal-to-conservative continuum. Widely used approaches involve the automatic processing of text documents produced by these actors, such as party manifestos (Laver et al., 2003; Slapin and Proksch, 2008) or legislative speeches (Lauderdale and Herzog, 2016). Other approaches involve human coding by experts (e.g., the Comparative Manifesto Project ) or crowd workers (Benoit et al., 2016). Yet other approaches rely on other inputs, such as roll call votes (Poole and Rosenthal, 1985), Twitter connections (Barberá, 2015), or campaign donations (Bonica, 2014).
@@
In this article, we propose a new approach to locate text documents in ideological and policy spaces using instruction-tuned Large Language Models (LLMs) and evaluate its performance. These models are LLMs optimized for dialogue use cases and are typically interacted with via chatbots such as ChatGPT. We focus on four scaling tasks using texts of different types, contexts, and lengths. First, we position political texts on the economic, social, and immigration policy dimensions based on party manifestos, a type of political text widely used in comparative politics research. Second, we position speeches by EU legislators, in 10 different languages, about a policy proposal on the ‘anti-subsidy’ to ‘pro-subsidy’ scale. Third, we position senators of the 117th US Congress on the left-right ideological spectrum based on the tweets they published during the Congress session. Fourth, we position individual tweets on the left-right spectrum that were published by US Representatives and Senators after the training cut-off date of the GPT-4 version used in these analyses (gpt-4-0613). [^1] The other LLMs with comparable performance analyzed in this research note were released in December 2023 or later and their training cutoff dates are not publicly available. This allows us to assess the performance of a modern LLM on prediction data we are certain were not part of the LLM pre-training data.
[…]
The scaling approach we propose is based on the direct query method introduced by Le Mens et al. (2023a), who used GPT-4 to measure the typicality of text documents in concepts. That article focused on several empirical settings, including the typicality of tweets in political parties (Democratic and Republican parties). We extend this previous work in four ways. First, we conduct a different task: we use LLMs for scaling text documents (party manifestos and legislative speeches) on particular policy dimensions rather than to measure their semantic similarity to a concept. Second, we use not only GPT-4 but also other LLMs, including open models (with architecture and weights available for download and use under permissive licenses). This overcomes the limitations regarding the future reproducibility of results produced with GPT-4 since the future availability of the version of the model used for research is not guaranteed by OpenAI. Third, we expand the text data and compare the validity of our estimates to those obtained using other scaling methods (e.g., the position of senators based on voting behavior). Fourth, we conduct performance evaluations not only on English texts but also in a multilingual context (EU legislative speeches).
@@
We know of only two recent studies that have employed instruction-tuned LLMs (ChatGPT) to estimate policy and ideological positions. Our approach to scaling differs from both. The first study used the LLM as a probabilistic text classifier to obtain the posterior probability that a sentence in a party manifesto is “Conservative” or “Liberal” and defined the position of the manifesto as the average (across sentences) of the difference between these probabilities. Our approach directly asks the LLM for the position of text on the focal dimension, and we find that it performs better. The second study directly asked ChatGPT to compare pairs of politicians on a particular dimension and used these pairwise comparisons to construct estimates of politician position on unidimensional scales (e.g., gun control support) (Wu et al., 2023). We ask LLMs to position political texts produced by political actors instead of asking them to position political actors based on their names. Our approach is thus applicable even for political actors who are not sufficiently talked about in the data used for pre-training the LLM to reliably answer questions about them.
@@ 2 Methods and data
Table 1 lists the LLMs we used for text scaling, including their open or closed status. These consist of a set of the most recent and largest LLMs available at the end of April 2024.222We also experimented with other open models (Mistral 7B, Meta LLama 2 chat 7B, 13B, 70B, Google Gemma 1.1 7B. These performed significantly worse than the open models included here (table A1 in the appendix).
To obtain a position estimate of a text document from an LLM, we submitted a prompt that contained a ‘user message’ that instructed it to return such an estimate. For example, to locate an excerpt of a party manifesto on the pro- to anti-immigration scale, we used:
You will be provided with a text from a party manifesto. $\ll$ Text from the party manifesto $\gg$ Where does this text stand on the ‘pro-immigration’ to ‘anti-immigration’ scale, in terms of immigration policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely pro-immigration’ and 100 means ‘Extremely anti-immigration’. To compute the score, start by first identifying the parts of the text that are about immigration policy. Then, compute the score based on these parts of text. If the text does not refer to immigration policy, return “NA”. You will only respond with a JSON object with the key Score. Do not provide explanations.
In all cases, we set the temperature parameter to 0, to ensure that the LLM would generate its response by selecting the most likely next token, and thus make the LLM responses as deterministic as possible (to ensure replicability). We also set the maximum number of tokens in the response to 20. This parameter does not affect the nature of the message returned by LLMs; it cuts the response down to 20 tokens if the LLM intended to generate a longer response. This ensures speed (token generation tends to be relatively slow) and limit costs (pay-per-use APIs charge per token submitted in the prompt and per token returned in the response). Finally, whenever this option was available, we set the response format to be a JSON object.
When the combination of text document and instructions resulted in a prompt longer than 4,096 tokens, we split the text document in chunks such that the prompts submitted to LLMs would not be longer than 4,096 tokens. The models we used can theoretically accept much longer prompts (up to 128K tokens for GPT-4 Turbo), but our tests revealed that submitting longer prompts leads to poorer performance. We obtained the position of each part of the text and took the weighted average of the scores for each part as the position of the text document (where the weight of a part was proportional to its number of tokens).
The appendix and the code shared on the companion OSF project provide the exact prompts and further details.
@@ 2.1 Data
2.1.1 British party manifestos and EU policy speeches in 10 languages
We obtained the texts, expert survey position estimates, and crowd-sourced position estimates from the replication package of Benoit et al. (2016). We scaled the British party manifestos analyzed in that paper on the same policy dimensions (economic and social policy for 18 manifestos, immigration policy for 8 manifestos). We also scaled (on the pro- to anti-subsidy dimension) the 36 speeches of a European Parliament debate on state subsidies. These were delivered in 10 different languages.
@@ 2.1.2 Senators of the 117th Congress
We obtained the list of senators from VoteView.com, their Twitter usernames, and downloaded the tweets they published during the Congress session through the Twitter API. Those who published fewer than 100 tweets during the Congress session were excluded from the analysis, resulting in a set of 98 senators.
We used the 900 tweets originally used in Le Mens et al. (2023a). In November 2023, we recruited 597 Prolific participants to each rate 30 tweets by answering the following question: “Where does this text stand on the ‘left’ to ‘right’ wing scale? If the text does not have political content, select ‘Not Applicable’.” Participants were not given any instructions as to what we mean by ‘left’ or ‘right.’ The crowdsourced position estimate of a tweet is the average of these ratings. All tweets, except one, received at least one position rating (different from ‘NA’), leading to a test data set of 899 tweets and their crowdsourced position estimates.
@@ 3.1 British party manifestos
Figure 1: Positioning British party manifestos on specific policy dimensions: Pearson correlations between the benchmark (expert survey placement analyzed in Benoit et al. (2016)) and either position estimates obtained with LLMs or fine-tuned BERT, or crowdsourced position estimates reported in Benoit et al. (2016). The numbers in parentheses indicate the number of documents for which the focal LLM returned a numeric position score.
@@
The position estimates obtained with the highest performing LLMs are very highly correlated with the Expert Survey Placements, more so than the crowdsourced position estimates (Figure 1). These results were obtained without providing any description of the policy dimension to the LLMs. Similar results were obtained by including such descriptions (Figure A5).
@@
To put these results into perspective, we used the crowdworkers’ position ratings collected by Benoit et al. (2016) to train a supervised probabilistic text classifier based on BERT, a transformer language model from an earlier generation than LLMs analyzed here (Devlin et al., 2018). We used the trained classifier to obtain position estimates of the manifestos’ sentences and, in turn, of the manifestos. This approach did not yield better results than those obtained with the best LLMs, although the latter were (most likely) not specifically trained to scale these party manifestos.
3.2 EU legislative speeches in 10 Languages
We obtained position estimates of each speech on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension, with instructions including background information.We used the original text of each speech, without translation or any other processing. Benoit et al. (2016) obtained 6 crowdsourced position estimates for each speech from crowdworkers coding the official translations in English, German, Greek, Italian, Polish, and Spanish. For each speech, we took the simple average of the six position estimates as a benchmark (Cronbach’s alpha reliability of .99). The correlation between this benchmark and the position estimates obtained with LLMs on the speeches in their original languages is .89 or higher for the highest performing LLMs (see upper right panel of Figure 2).
Refer to caption              Refer to caption Refer to caption
Figure 2: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension: Pearson correlations between the benchmark (crowdsourced position estimates reported in Benoit et al. (2016)). The numbers in parentheses indicate the number of documents for which a position was obtained.
For the best LLMs, the performance remains essentially the same when using official translations for both LLM and crowdsourcing, demonstrating that these LLMs can be used for scaling untranslated texts in multilingual contexts. As a performance benchmark, we report the pairwise correlations between crowdsourced estimates obtained from the translations of speeches in the different languages (upper left panel of Figure 2). These correlations are of similar magnitude.
3.3 Senators of the 117th US Congress
Figure 3: Positioning Senators of the 117th Congress on the left-right ideological spectrum based on their tweets ($N=98$). Each dot represents a senator (‘+’: Democrats, ‘x’: Republicans, ‘o’: others).
We queried the LLMs to obtain the positions, on the left-right ideological spectrum, of random samples of 100 tweets published by each senator during the 117th Congress session (January 3, 2021 to January 3, 2023). The position estimate of a senator is the average position of their tweets. Figure 3 shows that the resulting position estimates are highly correlated with those based on the roll call votes of the senators during that Congress session (Nokken and Poole, 2004), overall, and within-party. These position estimates are also highly correlated with those based on campaign funding (2020 CF scores, Bonica (2014)), but less so within party (Figure A14). Scaling the aggregated text of the 100 tweets in one or two prompts of up to 4,096 tokens produced similar results (Figure A15).
To compare these results with those obtained through approaches that do not require the submission of prompts to an LLM, we constructed position estimates for each tweet by computing their typicality in the Republican Party and their typicality in the Democratic Party by training probabilistic text classifiers (Le Mens et al., 2023b). We then defined the position of a tweet as the difference between these two typicalities. We used text classifiers based on fine-tuned BERT (the highest-performing approach in Le Mens et al. (2023b)), fine-tuned GloVe word embeddings, and a naive Bayes classifier based on word frequencies (TF-IDF) trained on 220K tweets by senators of the 116th and 117th US congresses. None of these approaches matches the best-performing LLMs, especially when it comes to capturing within-party differences.
Figure 4: Positioning tweets published by members of the US Congress after the training cut-off date of GPT-4 on the left-right ideological spectrum ($N=899$).
The LLMs did not return position estimates for all tweets but returned ‘NA’ for a subset of tweets, indicating that they judged that these tweets did not have enough political content to return a position estimate (see Appendix H for further discussion). The correlations between the position estimates produced by the best LLMs and crowdsourcing are very high, as shown in Figure 4. Position estimates reflect differences between-party and within-party. They are also higher than those produced by supervised classifiers trained on the approximately 1 million tweets published by members of the US Congress during the 116th and 117th Congress sessions.333 See  Appendix H.2 for an approach that asks LLMs to return typicality ratings. These results are all the more impressive considering that the comparison (crowdsourced position estimates) involves a benchmark of imperfect reliability (split-half correlation with Spearman-Brown correction; Overall: .95; Tweets by Democrats: .83; Tweets by Republicans: .86).
@@ 4 Discussion
These results demonstrate that “just asking” a modern instruction-tuned LLM for the position of text documents in ideological and policy spaces can produce valid position estimates, even with very short text documents such as tweets.
@@
This approach has the potential to greatly expand the scope of text analysis due to its accuracy, low cost, speed, ease of implementation and reproducibility (for open LLMs). Querying instruction-tuned LLMs is much less costly than human coding even when using pay-per-use APIs. Using GPT-4 to scale the 900 tweets analyzed in section 3.4 cost \$1.5 versus £1,626 for the crowdsourced position estimates; and running Meta Llama 3 on a laptop is free (though not as fast). Moreover, the accuracy of the position estimates produced with LLMs compares well to those produced by human coders. With the party manifestos, the position estimates of the best LLMs perform better than the crowdsourced position estimates, for all three policy dimensions. And with individual tweets, analyses reported in Appendix I show that the position estimates returned by the best LLMs are as accurate as the average of the ratings of about 4 or more independent human coders.
We have shown that some LLMs can reliably scale texts in their original languages, making the proposed approach particularly valuable for comparative research. But some LLMs with multilingual capabilities perform better in English than in other languages, probably because their pre-training data mainly consists of English-language texts. [^4] See for example, https://mistral.ai/news/mixtral-8x22b/. Systematic differences in accuracy across languages could lead to systematically different results in downstream analyses (e.g., a regression could lead to significant results in the English language setting, but not in the French language setting) simply due to the difference in the precision of the estimates obtained in the different languages. Addressing issues of differential measurement errors, bias, and their implications for downstream inference is necessary for LLMs to realize their full potential in comparative research.
@@
Which LLM should researchers who want to “ask an LLM” for the position of text documents in ideological and policy spaces choose? Five main considerations come into play: performance with respect to a relevant benchmark, cost, speed, data protection, and reproducibility. Until recently (Autumn 2023), commercial LLMs such as GPT-4 had a significant performance advantage on scaling tasks compared to the best open LLMs (Mistral 7B, Llama-2), but did not guarantee future reproducibility and involved sending data to OpenAI’s servers, creating potential data protection issues. This performance gap with the best open models has essentially vanished as of April 2024, at least in the comparative analyses we have performed for this research note. Therefore, we recommend that researchers use open LLMs such as Mixtral (8X22B) and Meta Llama 3 (70B) for text scaling tasks. They are cheaper and, most importantly, their architecture and weights are freely available for download, which guarantees a high level of data protection (if executed locally on a computer with enough memory) and the reproducibility of research results.
@@
“Asking an LLM might be all you need” to obtain accurate scaling results, but some caution is in order. Empirical validation and critical reflection on the scope of application of validation results are necessary steps for a scientifically valid use of LLMs for scaling and other measurement tasks in social and political science (for best practices, see the guidelines by Törnberg (2024)). The excellent performance of LLMs reported in this research note could tempt readers to use this approach in other domains while skipping the validation stage. But we advise them against doing so. LLMs are well-known for generating biased and unreliable results in some empirical settings. Until other researchers have shown that “just asking” a particular LLM provides accurate scaling results in a variety of empirical settings and focal latent dimensions, we cannot be certain about the breadth of settings in which LLMs perform well for scaling tasks and, a fortiori, other measurement or coding tasks. Moreover, evidence that an LLM performs well does not imply that others of the same generation will also perform well. We replicated some of the above analyses with Google Gemma 1.1. 7B (released in April 2024) and obtained much worse performance than with other recent LLMs. Until more is known about the range of domains in which LLMs perform well at scaling and other measurement tasks, case-by-case empirical validation remains essential.
5 Data Availability
Data and scripts have been deposited in OSF (https://osf.io/x2u5m/).
@@ 5 Data Availability
Prepared for "Scaling Political Texts with Large Language Models: Asking a Chatbot Might Be All You Need" by Gaël Le Mens and Aina Gallego
Appendix Appendix A Data
Appendix A.1 British party manifestos
The texts of the party manifestos used for scaling on the economic and social dimension were obtained from the file CSTA-APSR-master/DataCreated/master.sentences.csv which consisted of a table of sentences. We concatenated the sentences to construct the texts of the party manifestos.
CFjobsresultsImmigration/f354277_immigration1.csv.
Appendix A.2 EU legislative speeches
Appendix A.3 Senators of the 117th US Congress
Appendix A.4 Tweets published by members of the US Congress after the training cut-off date of GPT-4 – obtaining crowdsourced position estimates

### v3

# Positioning Political Texts with Large Language Models by Asking and Averaging

@@ Abstract
We use instruction-tuned Large Language Models (LLMs) like GPT-4, Llama 3, MiXtral, or Aya to position political texts within policy and ideological spaces. We ask an LLM where a tweet or a sentence of a political text stands on the focal dimension and take the average of the LLM responses to position political actors such as US Senators, or longer texts such as UK party manifestos or EU policy speeches given in 10 different languages. The correlations between the position estimates obtained with the best LLMs and benchmarks based on text coding by experts, crowdworkers, or roll call votes exceed .90. This approach is generally more accurate than the positions obtained with supervised classifiers trained on large amounts of research data. Using instruction-tuned LLMs to position texts in policy and ideological spaces is fast, cost-efficient, reliable, and reproducible (in the case of open LLMs) even if the texts are short and written in different languages. We conclude with cautionary notes about the need for empirical validation.
@@ 1 Introduction
Much research in the social and political sciences involves estimating the positions of actors, such as politicians or political parties, in latent ideological and policy spaces, such as the left-right or liberal-to-conservative continuum. Widely used approaches involve the automatic processing of text documents produced by these actors, such as party manifestos (Laver et al., 2003; Slapin and Proksch, 2008) or legislative speeches (Lauderdale and Herzog, 2016). Other approaches involve human coding by experts (Budge, 2001) or crowd workers (Benoit et al., 2016). Yet other approaches rely on other inputs, such as roll call votes (Poole and Rosenthal, 1985), Twitter connections (Barberá, 2015), or campaign donations (Bonica, 2014).
We propose a new approach to position text documents in ideological and policy spaces using instruction-tuned Large Language Models (LLMs) and evaluate its performance. These models are LLMs optimized for dialogue use cases and are typically interacted with via chatbots such as ChatGPT. We build on the direct query method introduced by Le Mens et al. (2023a) who measured the typicality of text documents in concepts by asking GPT-4 for typicality scores. We directly ask an LLM where a tweet or a sentence of a political text stands on the focal dimension and take the average of the LLM responses to obtain position estimates of longer texts or political actors.
@@
We focus on four scaling tasks using texts of different types, contexts, and lengths. First, we position individual tweets published by US Representatives and Senators by directly asking LLMs where these stand on the left-right ideological spectrum. Second, we position senators of the 117th US Congress on the same dimension by averaging the position estimates of a sample of the tweets they published during the Congress session. Third, we position party manifestos on the economic and social policy dimensions. We ask the LLMs for the positions of each sentence on these dimensions and average the LLM responses to obtain position estimates of the party manifestos. Fourth, we apply this approach to position speeches by EU legislators, in 10 different languages, about a policy proposal on the ‘anti-subsidy’ to ‘pro-subsidy’ scale. This allows us to explore the potential of this approach for comparative research with multilingual data.
@@
Several articles have shown that LLMs can produce text annotations (classifications in discrete categories such as relevant/irrelevant or a topic among a limited set of candidate topics) that are in good agreement with those produced by human coders (e.g., Gilardi et al., 2023; Törnberg, 2023; Ziems et al., 2023). However, little work has used LLMs to produce position estimates of political texts in ideological and policy spaces, which is the essence of text scaling, a core task in political science (see Benoit et al. (2020) for a discussion of the difference between the two tasks). We know of only two recent studies that rely on the text generation capabilities of LLMs to estimate policy and ideological positions. Our approach differs from both. The first study used GPT-3 as a probabilistic text classifier to obtain the posterior probability that a sentence in a party manifesto is “Conservative” or “Liberal” and defined the position of the manifesto as the average (across sentences) of the difference between these probabilities (Ornstein et al., 2024). Our approach directly asks the LLM for the position of text on the focal dimension, and we find that it performs better. [^1] The correlations between the Expert coding estimates and the positions produced by Ornstein et al. (2024) are .92.92.92.92 and .8.8.8.8. See Figures 3 and 4 for our results. The second study asked GPT 3.5 and Llama 2 to compare pairs of politicians on a particular dimension and used these pairwise comparisons to construct estimates of politician position on unidimensional scales (e.g., gun control support) (Wu et al., 2023). We ask LLMs to position political texts produced by political actors instead of asking them to position political actors based on their names. Our approach is thus applicable even to political actors about whom the LLM has little information.
@@ 2 Methods and data
2.1 Obtaining position estimates with LLMs
Table 1 lists the LLMs we used for text scaling, including their open or closed status. These consist of a set of the most recent and largest LLMs available at the end of May 2024.222See table A1 for analyses with other model such as Mistral, Gemma or Llama 2.
@@ 2.1 Obtaining position estimates with LLMs
To obtain a position estimate of a text with an LLM, we submitted a prompt that contained a ‘user message’ instructing it to return such an estimate. For example, to locate a tweet on the left-to-right-wing scale, we used:
You will be provided with the text of a tweet published by a member of the US Congress. Where does this text stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set the score to "NA". You will only respond with a JSON object with the key Score. Do not provide explanations. $\ll$ Text of the tweet $\gg$
In all cases, we set the temperature parameter to 0, to ensure that the LLM would generate its response by selecting the most likely next token, and thus make the LLM responses as deterministic as possible (to ensure replicability). We also set the maximum number of tokens in the response to 20. This parameter does not affect the nature of the message returned by LLMs; it cuts the response down to 20 tokens if the LLM intended to generate a longer response. This ensures speed (token generation tends to be relatively slow) and limits costs (pay-per-use APIs charge per token submitted in the prompt and per token returned in the response). Finally, whenever this option was available, we set the response format to be a JSON object.
To obtain the position of a party manifesto or a policy speech, we proceeded in a similar way with each sentence of the text documents. We then took the average of the positions of the sentences for which the LLM returned a numeric score, mimicking the approach used by Benoit et al. (2016) with human coders.
@@ 2.2.1 Tweets published by US Congress members after the training cut-off date of GPT-4
These data allow us to assess the performance of a modern LLM on prediction data we are certain were not part of the LLM pre-training data. We used the 900 tweets originally analyzed in Le Mens et al. (2023a). In November 2023, we recruited 597 Prolific participants to each rate 30 tweets by answering the following question: “Where does this text stand on the ‘left’ to ‘right’ wing scale? If the text does not have political content, select ‘Not Applicable’.” Participants were not given any instructions as to what we mean by ‘left’ or ‘right.’ The crowdsourced position estimate of a tweet is the average of these ratings. All tweets, except one, received at least one position rating (different from ‘NA’), leading to a test data set of 899 tweets and their crowdsourced position estimates. This measure is highly reliable overall and within-party (table 2).
Table 2: Reliability of the measures based on human ratings used as benchmark for assessing the performance of position estimates produced with LLMs.
@@ 2.2.2 Senators of the 117th Congress
We obtained the list of senators from VoteView.com, their Twitter usernames, and downloaded the tweets they published during the Congress session through the Twitter API. We used random samples of 100 tweets published by each senator during the 117th Congress session (January 3, 2021 to January 3, 2023). We excluded two senators who published fewer than 100 tweets during the Congress sessions. We use as a benchmark the first dimension Nokken-Poole period-specific DW-NOMINATE score, a well-established position estimate based on senators’ roll-call votes.
2.2.3 British party manifestos
We obtained the texts, expert coding estimates, and crowd coding estimates from the replication package of Benoit et al. (2016). We positioned the 18 British party manifestos on an economic policy dimension (from left- to right-wing) and on a social policy dimension (from conservative to liberal). We used as a benchmark the Expert Coding estimates that were constructed by Benoit et al. based on the sentence position estimates provided by a crowd of experts (political scientists). This measure is overall highly reliable and has varying levels of within-party reliability (table 2).
2.2.4 Multilingual setting: EU policy speeches in 10 languages
We also positioned the 36 speeches of a European Parliament debate on a policy proposal concerning state subsidies originally analyzed in Benoit et al. on the pro- to anti-subsidy dimension. These were delivered in 10 different languages by speakers who then voted for or against the proposal. Benoit et al. (2016) obtained 6 crowdsourced position estimates for each speech from crowdworkers coding the official translations in English, German, Greek, Italian, Polish, and Spanish. We took the simple average of the six crowd-coding estimates as a benchmark. This setting is challenging not only because of its multilingual nature but also because the speeches vary in style (e.g. technical, case-focused, rhetorical) and require knowledge of the debate context to be understood.
@@ 3.1 Tweets published by members of the US Congress after the training cut-off date of GPT-4
The LLMs returned ‘NA’ for a subset of tweets, indicating that they judged that these tweets did not have enough political content to return a position estimate (see E for further discussion). The correlations between the position estimates produced by the best LLMs and crowdsourcing are very high, as shown in Figure 1. Position estimates reflect differences between-party and within-party.
Figure 1: Positioning tweets published by members of the US Congress on the left-right ideological spectrum ($N=899$).
To compare these results with those obtained through approaches that do not require the submission of prompts to an LLM, we computed the typicality of each tweet in the Republican and the Democratic parties using probabilistic text classifiers and defined the position of a tweet as the difference between these two typicalities. The training data consist of approximately 1 million tweets published by members of the US Congress during the 116th and 117th Congress sessions.
We used text classifiers based on fine-tuned BERT (the highest performing approach in Le Mens et al. (2023b)), fine-tuned GloVe word embeddings and a naive Bayes classifier based on word frequencies (TF-IDF). [^3] See Appendix E.2 for an approach that asks LLMs to return typicality ratings. None of these approaches matches the best-performing LLMs, especially when it comes to capturing within-party differences.
3.2 Senators of the 117th US Congress
This setting differs from the previous one in that the benchmark positions are not based on human coding but on the voting behavior of the senators.
The position estimate of a senator is the average position of their tweets on the left-right ideological spectrum. Figure 2 shows that the resulting position estimates are highly correlated with those based on the roll call votes of the senators during that Congress session (Nokken and Poole, 2004), overall and within the party. These correlations are also higher than those produced by supervised classifiers used to obtain typicality measures of tweets in the two parties. The position estimates produced with LLMs are also highly correlated with those based on campaign funding (2020 CF scores, Bonica (2014)), although less so within-party (Figure A8).
Figure 2: Positioning Senators of the 117th Congress on the left-right ideological spectrum based on a random sample of 100 of their tweets ($N=98$). Each dot represents a senator (‘+’: Democrats, ‘x’: Republicans, ‘o’: others).
@@ 3.3 British party manifestos
Figure 3: Positioning British party manifestos on the Economic policy dimension (left to right wing scale). The numbers next to the dots indicate the years of the manifestos.
Figure 4: Positioning British party manifestos on the Social policy dimension (liberal to conservative scale). The numbers next to the dots indicate the years of the manifestos.
@@
The position estimates obtained with the highest performing LLMs are very highly correlated with the Expert Coding estimates, at a level comparable to the position estimates produced with crowd workers (Figures 3 and  4). This is the case not only overall, but also within political parties. These results were obtained without providing any description of the policy dimension to the LLMs. Similar results hold when including such descriptions (Figures A10 and  A11).
@@
We also trained a BERT-based supervised probabilistic text classifier (Devlin et al., 2018) using the crowdworkers’ ratings collected by Benoit et al. (2016), and used it to obtain position estimates of the manifestos’ sentences and, in turn, of the manifestos. This approach did not yield better results than those obtained with the best LLMs, although the latter were (most likely) not specifically trained to position these party manifestos.
3.4 Multilingual setting: EU policy speeches in 10 Languages
We obtained position estimates of the speeches on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension by submitting each sentence to the LLMs in its original language with instructions (in English) including background information on the context of the debate.
For the highest performing LLMs (GPT-4o and, to a lesser extent, GPT-4 Turbo, MiXtral 8X22B, Llama 3 70B, Aya 23 35B), the correlation between the benchmark and the position estimates obtained is high overall, and also when we separate the speeches by speakers who voted for and against the policy (Figure 5).
Results obtained with the translations of the speeches in the 6 languages used to obtain the crowd coding estimates show that the best models perform well across languages (Appendix H.2).
Figure 5: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension.
@@ 4 Discussion
These results demonstrate that “asking” modern instruction-tuned LLMs for the position of short texts in ideological spaces can produce valid position estimates. These can be used, in turn, to position political actors such as politicians, as we illustrated with US senators. We showed that asking LLMs for the positions of sentences in ideological and policy spaces and averaging the responses produces valid positions of party manifestos and policy speeches. The position estimates of the party manifestos produced with the best LLMs are as accurate as the crowdsourced position estimates. And with individual tweets, ancillary analyses show that the position estimates returned by the best LLMs are as accurate as the average of the ratings of about 4 or more independent human coders (Appendix E.3).
@@
This approach has the potential to expand the scope of text analysis due to its high accuracy, speed, ease of implementation, and reproducibility (for open LLMs). Moreover, querying instruction-tuned LLMs is much less costly than human coding even with the most expensive pay-per-use API (GPT-4): \$1.5 versus £1,626 for the 900 tweets analyzed in Section 3.1.
@@
Which LLM should researchers choose? Five main considerations come into play: accuracy with respect to a relevant benchmark, cost, speed, data protection, and reproducibility. If testing reveals no significant accuracy difference between open and closed LLMs, we recommend that researchers use open LLMs such as MiXtral (8X22B) and Llama 3 for text scaling tasks, at least in English. Their architecture and weights are freely available for download, which guarantees the reproducibility of research results and a high level of data protection (if executed locally or on a secure cloud machine).
For multilingual settings, it seems that GPT-4o (a proprietary model) has a marked advantage compared to the best open models at the time of writing, but some LLM designers are releasing open LLMs especially designed to perform well in multiple languages (e.g., the Aya series by Cohere). Some LLMs perform better in some languages than in others (Appendix H.2) and systematic accuracy differences across languages can bias the results of downstream econometric analyses. Developing approaches to deal with this differential measurement error would help realize the potential of LLMs in comparative research.
Another decision is whether to position long text documents in a single prompt or to split them into shorter parts, such as sentences. At this stage, we do not have a clear recommendation on this issue. Positioning party manifestos in a single prompt leads to lower performance than the sentence-by-sentence and averaging approach (Appendix G.2). In contrast, positioning senators by submitting their tweets in a single prompt did not cause significant performance degradation (Appendix F.2). Assessing where and when the single-prompt approach leads to performance degradation is an interesting avenue for future research.
When interpreting the results of the LLM-based approach for positioning political actors such as politicians or parties, it is important to remember that position estimates are based on the text documents submitted to the LLMs. Therefore, the validity of the resulting estimates is limited by the information contained in these texts. In the case of party positions, our approach resembles the approach of the Comparative Manifesto Project but differs from approaches that rely on surveys of experts about their perception of party positions, such as the Chapel Hill Expert Survey. This also implies that the results obtained with LLMs might differ from those obtained from expert surveys in the same way that those obtained by human coding can differ from those obtained with expert surveys because the inputs used to produce the position estimates differ.
@@
The high correlations between position estimates produced with LLMs and human coders reported in this research note could tempt readers to use this approach in other domains while skipping the validation stage. But we advise them against doing so. LLMs are well-known for generating biased and unreliable results in some empirical settings. Until other researchers have shown that asking (and averaging) a particular LLM provides accurate scaling results in a variety of empirical settings and focal latent dimensions, we cannot be sure about the breadth of settings in which LLMs perform well for scaling tasks and, a fortiori, other measurement or coding tasks. Until more is known about the range of domains in which LLMs perform well at scaling and other measurement tasks, case-by-case empirical validation remains essential.
@@
Prepared for "Positioning Political Texts with Large Language Models by Asking and Averaging" by Gaël Le Mens and Aina Gallego
Appendix A Data
A.1 Tweets published by members of the US Congress after the training cut-off date of GPT-4 – obtaining crowdsourced position estimates